Home‎ > ‎Reports‎ > ‎Crew098‎ > ‎

day03


Crew 98 Log Book for January 17, 2011

Photos of the day





Commander's Check-In Report

Virgiliu Pop Reporting

Crew Physical Status: Excellent

Time Departed/Returned from EVA: No EVA performed today.

Brief Narrative of Field Mission Results: N/A

EVA Data/Interpretations: N/A

Engineering/Hab Maintenance: As per the engineering report.

Report Transmission Schedule: CMDR check-in, CMDR report, Engineering, Photos

Plans for Tomorrow: Begin Sim

Inventory: Food boxes intended for our crew, and two big boxes for common use, were picked up from Hanksville.

Support Requested: Toilet needs to be fixed: clogged and no grey water available to flush it.

Miscellaneous: Last day out of sim.


Commander's Report

Virgiliu Pop Reporting

The last day out of sim begun at 9 AM with breakfast cooked by Betty and Iulia (scrambled eggs with mushrooms). Then Dragos and I headed to Hanksville in order to pick up the food allotted for our crew. We also retrieved the common food supplies. We then returned to the hab, where we all discussed in an open manner each member's strengths, weaknesses, and duties. Having people with different personalities in a confined environment is a great way of studying human behaviour, and we are sure that we will get along if we are open to compromise within the mission's limits.

Later on we used the last day out of sim for photo-ops back home - RoMars 2011 is followed by the Romanian public by the means of one of the main newspapers. Then the engineering team began its round. Their report is forthcoming by 8 PM. Some of us will head to Hanksville for the last time before beginning simulation, then we will eat dinner - cooked by the same Betty and Iulia - and we will set up the rover. Iulia and Betty will also catalogue the newly arrived food.

Tomorrow we will begin sim (including the food study), but before that we will clean the spacesuit helmets.

Engineering Report

Haritina Mogosanu Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Not used

Honey (Gas Generator) (Used? Yes/No): Yes, running all day

Battery Charge Level:
28.7 V green
33.7 V armor plate
40.3 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):
1 generator run all day ok


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.4

Gasoline Tank (Full/0.75/0.5/0.25/Empty):1 containers full x 5 gallons 2 half 1 a quarter.

Notes/Comments:


Oil: not checked we are not checking The oil and DG said to leave it on

Kitty Oil Quantity (# of Quarts in storage):

Honey Oil Quantity (# of Quarts in storage):

ATV Oil Quantity (# of Quarts in storage):on the engineering work bench 2 x 946 ml

Car Oil Quantity (# of Quarts in storage):where is it?


Water Status:

Outside Potable Water Tank Level (inches from bottom): 10

Trailer Potable Water Tank Level (inches from bottom): very low little water left

Hab Potable Water Tank Level (inches from bottom): 8

Potable Water Meter Reading: 409926

Notes/Comments:


GreenHab:

Condition of plants in Tank 1: plants dead please advise

Condition of plants in Tank 2: very poor only 4 plants left

Condition of plants in Tank 3: poor

Greenhab Notes/Comments: plants are very few and have some leaves dead.


Observatory:

Telescope Used? (Yes/no) yes

Observatory Notes: dome engine felt off the rack and needs re-bolting. Contacted both Peter Detterline and DG.

(Note from Mission Support: the dome engine was removed a year ago due to warping - so the dome should be opened by hand.)


Transportation:

Hab Car (NOT RENTAL)
Used (yes/no):Yes
Oil Checked (yes/no):No
Tire Status:not checked
Odometer at end of day:-123152
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: filled
Tire Status:not checked
Comments:

ATV 2 (Spirit):y
Used: y/n y
Oil Checked: y/n n
Fuel Consumed:replaced
Tire Status:not checked
Comments:

ATV 3 (Viking I):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status: not checked
Comments:

ATV 4 (Viking II):
Used: y/n y
Oil Checked: y/n n
Fuel Consumed: replaced
Tire Status:not checked
Comments:

Took the ATVs for a check drive around the habitat not on the road.


Heating and Ventilation:
Thermostat setting upstairs (Fahrenheit): 65
Actual temperature upstairs (Fahrenheit): 78

Thermostat setting downstairs (Fahrenheit): 55
Actual temperature downstairs (Fahrenheit): 75


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: all ok


General Engineering Notes/Comments: can we organize the junk yard? (we were thinking of tidying it up)

Journalist Report

Dragos Bratasanu Reporting

Today was an extremely busy day for all the crew members. After the morning engineering round, I went with the commander to the village to get the food supplies and ran back to the hab.

Iulia and Beatrice worked all day on setting up the rover for the following days and went through geology books, and myself and Haritina set up the observatory. There was a lot of work to reach the start and we will also continue tomorrow because we didn't finish.

The evening ended wonderfully. We took the official pictures of the crew, as attached. Hope you enjoy them!